OpenStudy (anonymous):

A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. How much of the 70% solution is used?

OpenStudy (owlfred):

Let's think in terms of pure acid. I love pure acid. 50 liters of a 40% acid solution contains \(50 \times \frac{40}{100} = 20 ~l\) of pure acid. That's a good start. Now, let the volume of \(20\%\) acid solution be \(x\) and the volume of \(70\%\) acid solution be \(y\). First of all, mixing those two solutions gives you 50 liters of a solution, meaning that \(x+ y = 50\). Secondly, let's again think in terms of pure acid. If I have \(x\) liters of 20% acid solution, then I will have \(\frac{20}{100}x\) liters of pure acid. Also, if I have \(y\) liters of a 70% acid solution, then I have \(\frac{70}{100}y \) liters of pure acid. When I mix those two, I get 20 liters of pure acid. In other words, \(\text{pure acid from the first mixture + pure acid from the second mixture}\) is equal to \(\text{20 liters of pure acid}\). Our second equation would then be \(\frac{2}{10}x + \frac{7}{10}y = 20\).